OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)? A. 2 B. -4/5 C. 5/4 D. -2

OpenStudy (calculusxy):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 } = \frac{ 1 - (-5) }{ 4 - 1 } = \frac{ 6 }{ 3 } = 2\]

OpenStudy (calculusxy):

there various formulas for the slope which all mean the same thing: \[\frac{ rise }{ run } = \frac{ y_2 - y_1 }{ x_2 - x_1 } = \frac{ \Delta y }{ \Delta x }\]

OpenStudy (calculusxy):

you can user either one of them to find the slope

OpenStudy (calculusxy):

*use